# A Comprehensive Survey on Continual Learning in Generative Models

### v3

@@ 6.2 Architecture-based Approach
C-LoRA [132] introduces a continual concept learning setting within the LDM framework, where each user-defined concept is assigned a dedicated LoRA module injected into the self-attention keys and values. During training, each concept is represented by a unique placeholder token (e.g., [$\mathcal{V}^{*}$]) inserted into the prompt to associate the visual concept with a textual identifier. To mitigate catastrophic forgetting, C-LoRA freezes the LoRA parameters associated with previously learned concepts during new concept learning. Furthermore, it incorporates a self-regularization loss to penalize updates to weight regions that were modified by earlier concepts, thus preserving prior knowledge.
6.3 Regularization-based Approach
@@
MuseumMaker [133] aims to retain the knowledge of past learned styles while incorporating new customized styles. To mitigate catastrophic overfitting and maintain the integrity of original concepts in diffusion models, it introduces a style distillation loss that captures disentangled style representations. Furthermore, it proposes a dual regularization scheme for LoRA training, which constrains model optimization from both the weight and feature perspectives to achieve balanced knowledge integration. LFS-Diffusion [134] proposes a data-free knowledge distillation strategy applied across diffusion time steps during training, aiming to constrain the model from overfitting to current personalized concepts and thereby preserve the original generative capabilities. During inference, an in-context generation mechanism is introduced to mitigate forgetting of previously learned knowledge and ensure better retention of past concepts. CIDM [135] introduces a concept consolidation loss that employs learnable layer-wise concept tokens and an orthogonal subspace regularizer to disentangle task-specific and task-shared knowledge. During inference, an elastic weight aggregation module merges low-rank weights of previously learned concepts based on their importance. To alleviate concept neglect in multi-concept composition, CIDM further proposes a context-controllable synthesis strategy that enhances regional feature expressiveness via layer-wise textual embeddings and guides generation through region-aware noise estimation. ConceptGuard [136] introduces a set of lightweight yet effective strategies to mitigate forgetting in continual concept learning for diffusion models. It employs shift embeddings for dynamic concept updates, concept-binding prompts to disentangle and compose concepts over time, and memory preservation regularization to slow down harmful parameter drift. Additionally, a priority queue adaptively manages concept replay based on their importance and temporal order. Jha et al. [137] utilize diffusion classifier scores as implicit class-conditional density estimators. This method simultaneously regularizes both parameter-space and function-space consolidation, enabling effective adaptation to evolving user concepts.
6.4 Replay-based Approach
@@
Since each stage in the continual personalized diffusion setting typically involves only a limited number of training samples (e.g., 3-5 images), a feature replay strategy is commonly adopted. Specifically, $\text{L}^{2}\text{DM}$ [138] addresses the continual concept learning challenge by introducing the Rainbow-Memory Bank Strategy, which consists of a short-term memory bank for encoding current task features, a long-term memory bank for preserving historical knowledge, and a scoring function to balance knowledge retention and transfer. To further mitigate the forgetting of personalized details, $\text{L}^{2}\text{DM}$ proposes Elastic Concept Distillation, which enhances model performance by distilling knowledge between the current model and the previous task model, using features retrieved from the short-term memory bank as input to both.
6.5 Discussions
@@ 7 Future Directions
Efficient Continual Learning Mechanisms. Current mainstream continual learning approaches often rely on model expansion to retain task-specific knowledge, which leads to substantial storage and computational overhead during deployment. These challenges are particularly pronounced for large-scale generative models, where the large number of parameters causes latency and resource limitations. Consequently, there is increasing interest in enhancing efficiency through techniques such as model compression [254, 255, 256], pruning [257, 258, 259, 260, 261, 262], and quantization [263, 264, 265]. In addition to structural optimization, recent progress in dataset compression and distillation [266, 267, 268, 269] highlights the potential of compressing replay data as a promising direction. A key idea is to encode as much task-relevant information as possible from previous data or features into a small number of representative tokens. This strategy not only helps mitigate catastrophic forgetting but also reduces the risk of data privacy leakage by avoiding the storage of raw data. Overall, integrating continual learning with efficient model and data compression techniques provides a compelling path toward improving the practicality and scalability of generative models in real-world applications.
@@
Learning Paradigm for Continual Learning. While conventional continual learning methods predominantly rely on supervised fine-tuning (SFT) for incremental knowledge acquisition, emerging reinforcement learning (RL)-based training paradigms, particularly exemplified by GRPO [270], have demonstrated stronger generalization capabilities across a variety of model architectures. This performance advantage has been consistently observed not only in large language models [271, 272, 273] but also in multimodal large language models [274, 275, 276, 277] and vision-language action models [278, 279], suggesting that RL-based methods may serve as a more robust alternative to traditional SFT strategies in continual learning settings. Recent studies [280] have revealed a fundamental difference between the two paradigms: SFT tends to overfit specific training distributions through pattern memorization, whereas RL emphasizes structural comprehension of task objectives, enabling better generalization to novel scenarios. This distinction provides strong motivation for exploring continual learning within an RL framework, which may naturally support long-term adaptation, task abstraction, and reinforcement-driven retention strategies.
@@
Self-generation for Continual Learning. A core challenge in continual learning lies in the inaccessibility of previous task data. In discriminative models, a widely adopted solution involves training auxiliary generative models (e.g.,, GANs [281] or diffusion models) to synthesize pseudo-samples of prior tasks, thereby mitigating catastrophic forgetting [282, 283, 284, 285]. While effective, this approach introduces significant computational overhead and may suffer from a distributional mismatch between generated and original data. In contrast, generative models offer a more integrated solution by leveraging their inherent capacity to reconstruct and regenerate data. These models can autonomously replay past task distributions during the learning of new tasks, without the need for separate generative modules. This capability offers two major benefits: reducing additional architectural complexity and achieving better alignment with the original data distribution. As a result, self-generation in generative models presents a promising direction for building more efficient, scalable, and distributionally consistent continual learning systems.
@@
Continual Learning in Large Scale Models. Most existing studies on continual learning for generative models have focused on small- to medium-scale architectures (e.g., models with up to 7B and 13B parameters). However, emerging evidence [286, 287] suggests that continual learning behaviors can differ substantially as the model scale increases. This observation highlights two critical research directions. First, there is a pressing need to establish specialized evaluation benchmarks that mitigate information leakage risks [288], particularly given the large and often proprietary datasets used to train large-scale models. Second, it is essential to explore parameter-efficient knowledge integration strategies—such as Chain-of-Thought reasoning (CoT)[166, 289], In-Context Learning (ICL)[290, 291, 292], Agents [293, 294, 295], and Retrieval-Augmented Generation (RAG) [296, 297]—as conventional fine-tuning becomes increasingly infeasible for models with extensive parameter counts.
@@
Continual Learning for Multimodal Generation Beyond Vision and Language. While most existing research in continual generative learning focuses on vision-language tasks, real-world multimodal systems are increasingly incorporating richer modalities such as audio [298], 3D scenes [299], and motion data [300]. In such settings, modality-continual learning becomes a critical yet underexplored challenge. For example, a pre-trained MLLM may be required to incorporate speech input or generate spoken responses. Without appropriate continual learning mechanisms, learning the new audio modality could result in severe forgetting of previously acquired vision-language capabilities. Similarly, in VLA models deployed in robotics or embodied agents, the integration of new sensor modalities (e.g., depth and haptic signals) necessitates stable expansion of generative capacity without overwriting existing knowledge. These scenarios highlight the need for continual learning methods capable of integrating new modalities while preserving alignment and generative quality across previously learned ones.
@@
Toward Unified Optimization Frameworks for Continual Learning in Generative Models. Although generative models vary widely in data modalities and task objectives, increasing architectural convergence is fostering the development of unified continual learning frameworks. For example, both multimodal large language models and vision-language action models frequently rely on large language models to enable cross-modal reasoning. Furthermore, recent attempts to integrate diffusion models with MLLMs [301, 302, 303] underscore a broader trend toward unifying generation and understanding within generative AI. These shared characteristics suggest a promising direction: designing continual learning strategies that generalize across generative model families. Unifying optimization across modalities and architectures may mark a critical step toward deploying generative AI in complex, dynamic real-world scenarios.
8 Conclusion
@@
In this survey, we have reviewed recent advances in continual learning for generative models, providing a systematic overview across four major categories: Large Language Models, Multimodal Large Language Models, Vision-Language Action Models, and Diffusion Models. We further unify these models by identifying shared methodological dimensions that underpin their continual learning paradigms. Specifically, we formalize the continual learning problem in each domain, introduce relevant datasets and model architectures, and categorize representative approaches within each subfield. We hope that this comprehensive and in-depth survey, along with the accompanying real-time updated GitHub repository, will not only consolidate the current understanding of continual learning in generative modeling but also serve as a valuable reference and catalyst for future research.

### v4

@@ 6.2. Architecture-based Approach
C-LoRA (Smith et al., 2023) introduces a continual concept learning setting within the LDM framework, where each user-defined concept is assigned a dedicated LoRA module injected into the self-attention keys and values. During training, each concept is represented by a unique placeholder token (e.g., $\mathcal{V}^{*}$]) inserted into the prompt to associate the visual concept with a textual identifier. To mitigate catastrophic forgetting, C-LoRA freezes the LoRA parameters associated with previously learned concepts during new concept learning. Furthermore, it incorporates a self-regularization loss to penalize updates to weight regions that were modified by earlier concepts, thus preserving prior knowledge.
6.3. Regularization-based Approach
@@
MuseumMaker (Liu et al., 2024b) retains previously learned styles while integrating new custom ones. To prevent overfitting and preserve original diffusion concepts, it introduces a style distillation loss that captures disentangled style representations. It also proposes a dual regularization scheme for LoRA training, constraining optimization from both weight and feature perspectives to balance knowledge integration. LFS-Diffusion (Song et al., 2024) adopts a data-free distillation strategy across diffusion time steps to curb overfitting to new concepts and preserve generative capacity. In inference, an in-context generation mechanism enhances the retention of prior knowledge. CIDM (Dong et al., 2024) proposes a concept consolidation loss using learnable layer-wise concept tokens and an orthogonal subspace regularizer to separate task-specific and shared knowledge. It further introduces elastic weight aggregation for merging concept-specific LoRA weights and a context-controllable synthesis strategy that enhances regional features via layer-wise text embeddings and region-aware noise. ConceptGuard (Guo and Jin, 2025) mitigates forgetting through shift embeddings for dynamic updates, concept-binding prompts for composition, and memory preservation regularization to limit parameter drift. A priority queue manages concept replay based on importance and recency. Jha et al. (Jha et al., 2025) leverage diffusion classifier scores as implicit class-conditional density estimators to guide continual adaptation. By regularizing both parameter and function spaces, their method enables diffusion models to adapt to evolving user concepts while minimizing forgetting of prior knowledge.
6.4. Replay-based Approach
@@
Since each stage in the continual personalized diffusion setting typically involves only a limited number of training samples (e.g., 3-5 images), a feature replay strategy is commonly adopted. Specifically, $\text{L}^{2}\text{DM}$ (Sun et al., 2024) addresses the continual concept learning challenge by introducing the Rainbow-Memory Bank Strategy, which consists of a short-term memory bank for encoding current task features, a long-term memory bank for preserving historical knowledge, and a scoring function to balance knowledge retention and transfer. To further mitigate the forgetting of personalized details, $\text{L}^{2}\text{DM}$ proposes Elastic Concept Distillation, which enhances model performance by distilling knowledge between the current model and the previous task model, using features retrieved from the short-term memory bank as input to both.
6.5. Discussions
@@ 7. Future Directions
Efficient Continual Learning Mechanisms. Current mainstream continual learning approaches often rely on model expansion to retain task-specific knowledge, which leads to substantial storage and computational overhead during deployment. These challenges are particularly pronounced for large-scale generative models, where the large number of parameters causes latency and resource limitations. Consequently, there is increasing interest in enhancing efficiency through techniques such as model compression (Zhu et al., 2024a; Li et al., 2024; Chen et al., 2025b), pruning (Cheng et al., 2024; Wu et al., 2023b; An et al., 2024; Zeng et al., 2025b; Castells et al., 2024; Kong et al., 2025), and quantization (Liu et al., 2024c; Zhao et al., 2024a; Lin et al., 2024). In addition to structural optimization, recent progress in dataset compression and distillation (Wang et al., 2018; Liu et al., 2022; Ma et al., 2025; Wang et al., 2025b) highlights the potential of compressing replay data as a promising direction. A key idea is to encode as much task-relevant information as possible from previous data or features into a small number of representative tokens. This strategy not only helps mitigate catastrophic forgetting but also reduces the risk of data privacy leakage by avoiding the storage of raw data. Overall, integrating continual learning with efficient model and data compression techniques provides a compelling path toward improving the practicality and scalability of generative models in real-world applications.
@@
Learning Paradigm for Continual Learning. While conventional continual learning methods predominantly rely on supervised fine-tuning (SFT) for incremental knowledge acquisition, emerging reinforcement learning (RL)-based training paradigms, particularly exemplified by GRPO (Shao et al., 2024), have demonstrated stronger generalization capabilities across a variety of model architectures. This performance advantage has been consistently observed not only in large language models (Guo et al., 2025a; Li et al., 2025b; Yu et al., 2025b) but also in multimodal large language models (Liu et al., 2025a; Huang et al., 2025a; Shen et al., 2025; Peng et al., 2025) and vision-language-action models (Guo et al., 2025d; Zhang et al., 2024c), suggesting that RL-based methods may serve as a more robust alternative to traditional SFT strategies in continual learning settings. Recent studies (Chu et al., 2025) have revealed a fundamental difference between the two paradigms: SFT tends to overfit specific training distributions through pattern memorization, whereas RL emphasizes structural comprehension of task objectives, enabling better generalization to novel scenarios. This distinction provides strong motivation for exploring continual learning within an RL framework, which may naturally support long-term adaptation, task abstraction, and reinforcement-driven retention strategies.
@@
Self-generation for Continual Learning. A core challenge in continual learning lies in the inaccessibility of previous task data. In discriminative models, a widely adopted solution involves training auxiliary generative models (e.g.,, GANs (Goodfellow et al., 2014) or diffusion models) to synthesize pseudo-samples of prior tasks, thereby mitigating catastrophic forgetting (Shin et al., 2017; Wu et al., 2018; Cong et al., 2020; Meng et al., 2024). While effective, this approach introduces significant computational overhead and may suffer from a distributional mismatch between generated and original data. In contrast, generative models offer a more integrated solution by leveraging their inherent capacity to reconstruct and regenerate data. These models can autonomously replay past task distributions during the learning of new tasks, without the need for separate generative modules. This capability offers two major benefits: reducing additional architectural complexity and achieving better alignment with the original data distribution. As a result, self-generation in generative models presents a promising direction for building more efficient, scalable, and distributionally consistent continual learning systems.
@@
Continual Learning in Large Scale Models. Most existing studies on continual learning for generative models have focused on small- to medium-scale architectures (e.g., models with up to 7B and 13B parameters). However, emerging evidence (Bi et al., 2024; Zhang et al., 2024b) suggests that continual learning behaviors can differ substantially as the model scale increases. This observation highlights two critical research directions. First, there is a pressing need to establish specialized evaluation benchmarks that mitigate information leakage risks (Kim et al., 2023), particularly given the large and often proprietary datasets used to train large-scale models. Second, it is essential to explore parameter-efficient knowledge integration strategies—such as Chain-of-Thought reasoning (CoT)(Wei et al., 2022; Zhang et al., 2023e), In-Context Learning (ICL)(Dong et al., 2022; Momeni et al., 2024; Xu et al., 2024), Agents (Xi et al., 2025; Xie et al., 2024a; Zheng et al., 2025a), and Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023)—as conventional fine-tuning becomes increasingly infeasible for models with extensive parameter counts.
@@
Continual Learning for Multimodal Generation Beyond Vision and Language. While most existing research in continual generative learning focuses on vision-language tasks, real-world multimodal systems are increasingly incorporating richer modalities such as audio (Zhang et al., 2023a), 3D scenes (Hong et al., 2023), and motion data (Jiang et al., 2023). In such settings, modality-continual learning becomes a critical yet underexplored challenge. For example, a pre-trained MLLM may be required to incorporate speech input or generate spoken responses. Without appropriate continual learning mechanisms, learning the new audio modality could result in severe forgetting of previously acquired vision-language capabilities. Similarly, in VLA models deployed in robotics or embodied agents, the integration of new sensor modalities (e.g., depth and haptic signals) necessitates stable expansion of generative capacity without overwriting existing knowledge. These scenarios highlight the need for continual learning methods capable of integrating new modalities while preserving alignment and generative quality across previously learned ones.
@@
Toward Unified Optimization Frameworks for Continual Learning in Generative Models. Although generative models vary widely in data modalities and task objectives, increasing architectural convergence is fostering the development of unified continual learning frameworks. For example, both multimodal large language models and vision-language-action models frequently rely on large language models to enable cross-modal reasoning. Furthermore, recent attempts to integrate diffusion models with MLLMs (Xie et al., 2024b; Ge et al., 2024; Wu et al., 2024a) underscore a broader trend toward unifying generation and understanding within generative AI. These shared characteristics suggest a promising direction: designing continual learning strategies that generalize across generative model families. Unifying optimization across modalities and architectures may mark a critical step toward deploying generative AI in complex, dynamic real-world scenarios.
8. Conclusion
@@
In this survey, we have reviewed recent advances in continual learning for generative models, providing a systematic overview across four major categories: Large Language Models, Multimodal Large Language Models, Vision-Language-Action Models, and Diffusion Models. We further unify these models by identifying shared methodological dimensions that underpin their continual learning paradigms. Specifically, we formalize the continual learning problem in each domain, introduce relevant datasets and model architectures, and categorize representative approaches within each subfield. We hope that this comprehensive and in-depth survey, along with the accompanying real-time updated GitHub repository, will not only consolidate the current understanding of continual learning in generative modeling but also serve as a valuable reference and catalyst for future research.
Acknowledgements.
This work was supported by the National Science and Technology Major Project (2022ZD0116500), National Natural Science Foundation of China (62222609, 62320106010), CAS Project for Young Scientists in Basic Research (YSBR-083), Major Science and Technology Plan Project on the Future Industry Fields of Xiamen City (3502Z20241027), Unveiling and Leading Projects of Xiamen (3502Z20241011) and the InnoHK program.